kristen00001:

. Simplify the expression -9x+3(4-x)

umm:

Given:\[-9x+3(4-x)\]Distribute\[=−9x+(3)(4)+(3)(−x)\]\[=−9x+12+−3x\]Now you combine like terms\[=−9x+12+−3x\]\[=(−9x+−3x)+(12)\]\[=−12x+12\]Therefore your answer is:\[=−12x+12\]

minustempo:

So first you need to distribute the 3(4-x). Do you know the distributive property?

dukedadon:

lit

@umm wrote:
Given:\[-9x+3(4-x)\]Distribute\[=−9x+(3)(4)+(3)(−x)\]\[=−9x+12+−3x\]Now you combine like terms\[=−9x+12+−3x\]\[=(−9x+−3x)+(12)\]\[=−12x+12\]Therefore your answer is:\[=−12x+12\]
and your source is?

minustempo:

source is the brain

umm:

My brain. (: This is simple math.

dukedadon:

@umm wrote:
My brain. (: This is simple math.
hmm .

kristen00001:

uh-thanks ig

ILOVESPAGHETTI:

@umm wrote:
My brain. (: This is simple math.
Thank you for doing that. I was not ready to type all that out

dukedadon:

thats crazy because i just looked on google and .. just randomly, i saw your answer in the same exact form

dukedadon:

care to explain ssl?

umm:

You do realize this is simple math, no?

minustempo:

there's only 1 way to do it

ILOVESPAGHETTI:

@dukedadon wrote:
care to explain ssl?
You always need to use the exact same steps for math

minustempo:

@ilovespaghetti wrote:
@dukedadon wrote:
care to explain ssl?
You always need to use the exact same steps for math
not necessarily, only for basic math

dukedadon:

@umm wrote:
You do realize this is simple math, no?
if it is so simple, then why does google have the exact answer as yours?

umm:

Closed: Question has been answered, move along. This can be discussed in chat(s).

kristen00001:

why yall still talkin abt this question????

dukedadon:

oh now its a closed question ..

umm:

@dukedadon wrote:
if it is so simple, then why does google have the exact answer as yours?
Because, there's only one correct answer to this.

kristen00001:

uh yea thanks